How to apply telescopic series in series involving sine of angles

Grade:11

3 Answers

Akshay
8 years ago
suppose your seies is : sin x + sin 2x + sin 3x..........sin nx,
multiply and divide by sin x/2,
use sin x * sin y = ½ * (cos(x-y) – cos(x+y) ),
sin x * sinx/2 = 0.5 * cos(x/2) – 0.5 * cos(1.5x),
sin 2x * sin x/2 = 0.5 * cos(1.5 x) – 0.5*cos(2.5 x),
sin 3x * sin x/2 = 0.5 * cos(2.5 x) – 0.5*cos(3.5 x),
and so on,      adding them all of the terms will get cancelled out.
Akshay
8 years ago
or if you want to do it using exponentials,
write sin ax =1/2 * (eiax-e-iax),
u will get two geomtric series, one for +ve powers of exp, other for negative powers. apply sum of gm formula 
manu
8 years ago
se sin x * sin y = ½ * (cos(x-y) – cos(x+y) ),
sin x * sinx/2 = 0.5 * cos(x/2) – 0.5 * cos(1.5x),
sin 2x * sin x/2 = 0.5 * cos(1.5 x) – 0.5*cos(2.5 x),
sin 3x * sin x/2 = 0.5 * cos(2.5 x) – 0.5*cos(3.5 x),